OpenStudy (anonymous):

Find the logarithm log­(1/4) 64. A. -2 B. -3 C. -4 D. -16

OpenStudy (anonymous):

I got -3

OpenStudy (anonymous):

\[\log_{\frac{1}{4}}(64)\] solve\[\left(\frac{1}{4}\right)^x=64\]

OpenStudy (anonymous):

-3 is right

OpenStudy (anonymous):

okauy:)

OpenStudy (anonymous):

-3 is correct. The negative sign flips 1/4 into 4/1 and 4^3 = 64